# Learning Value of Information towards Joint Communication and Control in 6G V2X

### v1

@@ VII-C Optimization Objective
Case 9.
In Case 8, the optimization objective of the communication agents $i-1\in\mathcal{L}$ is to maximize the sum V2I throughtput while minimizing the performance degradation of the followers $i$ due to non-ideal communications.
@@
Similarly, let $G_{m,(k,t)}$ denote the instantaneous channel gain of the V2I link $m$; $G_{i-1,B,m,(k,t)}$ the interference channel gain from V2V link $i-1$ transmitter at predecessor $i-1$ to V2I link $m$ receiver at the BS; $G_{B,i-1,m,(k,t)}$ the interference channel gain from V2I link $m$ transmitter at the BS to V2V link $i-1$ receiver at follower $i$; and $G_{j-1,i-1,m,(k,t)}$ the interference channel gain from the V2V link $j-1$ transmitter at predecessor $j-1$ to the V2V link $i-1$ receiver at follower $i$ over the sub-channel $m$.
@@
$\gamma_{m,(k,t)}=\frac{P^{\rm I}_{m}G_{m,(k,t)}}{\sigma^{2}+\sum\limits_{i-1\in\mathcal{L}}\phi_{i-1,k}\theta_{i-1,m,(k,t)}P^{\rm V}_{i-1,m,(k,t)}G_{i-1,B,m,(k,t)}},$ (36)
@@
respectively. $\sigma^{2}$ is the power of channel noise which satisfies the independent Gaussian distribution with a zero mean value. $I_{i-1,m,(k,t)}$ is the total interference power received by V2V link $i-1$ over sub-channel $m$, where
@@
Each predecessor $i-1\in\mathcal{L}$ has a buffer to store its CAM. Let $q^{\rm CAM}_{i-1,(k,t)}$ denote the queue length of predecessor $i-1$, in terms of the number of CAMs, at communication interval $(k,t)$. At the beginning of the first communication interval $(k,0)$ of each control interval $k\in\mathcal{K}$, predecessor $i-1$ samples the driving status for control interval $k$ and buffers the generated CAM. At each communication interval $(k,t),t\in\mathcal{T}$, within control interval $k$, the queue length $q^{\rm CAM}_{i-1,(k,t)}$ is decreased by $\Delta t\times r^{\rm CAM}_{i-1,(k,t)}$, which is the number of CAM transmitted during the communication interval. Since any CAM that is not fully transmitted by the end of the last communication interval $(k,T-1)$ of control interval $k$ will be discarded, at the start of each new control interval $k\in\mathcal{K}$, specifically at communication interval $(k,0)$, $q^{\rm CAM}_{i-1,(k,0)}$ is set to $1$. Since the value of $q^{\rm CAM}_{i-1,(k,0)}$ changes abruptly at the beginning of each control interval $k$, we use $q^{\rm CAM}_{i-1,(k,T)}$ to denote the queue length by the end of the last communication interval $(k,T-1)$ of control interval $k$. Note that $q^{\rm CAM}_{i-1,(k,T)}\neq q^{\rm CAM}_{i,(k+1,0)}$ since the queue length is discontinuous at the control interval boundaries.
@@
$\left\{\begin{array}[]{ll}1,&\mathrm{if}\quad t=0\\ \max\left[0,q^{\rm CAM}_{i-1,(k,t-1)}-\Delta t\times r^{\rm CAM}_{i-1,(k,t-1)}\right],&\mathrm{if}\quad 0<t\leq T\\ \end{array}\right..$ (43)
For each follower $i$, if a CAM is fully delivered to it by the end of control interval $k$, the receiver will replace its old CAM with the newly received CAM at the beginning of control interval $k+1$. If a CAM is not fully delivered by the end of the control interval, the receiver will continue to use the last fully received CAM, as it cannot interpret an incomplete CAM. When the most recent available CAM of follower $i$ at the beginning of control interval $k$ corresponds to the CAM sampled at the predecessor $i-1$ in control interval $k-\tau_{i,k}$, we refer to $\tau_{i,k}$ as the observation delay at V2V receiver $i$, which is derived from the queue length as
@@
Notably, each follower $i$ makes decisions based on the augmented state of a Random Delay SSDP as given in Case 5.
@@
where $\gamma$ is the RRM reward discount factor. The first term corresponds to the discounted sum throughput of all V2I links $m\in\mathcal{M}$ over all the communication intervals of a control episode, and the second term corresponds to the EVoMI $\Xi_{i}^{\rm M}$ or EVoII $\Xi_{i}^{\rm I}$ of follower $i$ defined in Case 6. The weight factors $\kappa_{1}$ and $\kappa_{2}$ indicate the relative importance of minimizing control performance degradation versus maximizing the V2I throughput.
@@
where the expectation is taken with respect to the probability distribution of the state-action trajectories when the RRM policy is $\pi^{\rm CM}$ and vehicle control policy is $\pi^{\mathrm{inf}}$.
@@ VII-D VoI-oriented Optimization for Vehicular Communication
When addressing the problem of ”When and What to Communicate”, different approaches might be adopted depending on the decision-making time scale, i.e.,
  • Across Multiple Control Intervals: When decisions are made statically or semi-statically across multiple control intervals, we can evaluate the $J^{\mathrm{CM}}$ for different decisions and choose the decision that maximizes $J^{\mathrm{CM}}$. To evaluate $J^{\mathrm{CM}}$, the EVoI can be estimated as outlined in Section V, and $J^{\mathrm{CM\_trad}}$ can be evaluated using theoretical tools such as queuing theory or through simulation.
  • Every Control Interval: When decisions are made dynamically per control interval, we can formulate an SSDP to optimize the communication decisions made at each control interval. Specifically, the reward function of the SSDP can be written as
    $r^{\rm CM}_{(k,t)}=\kappa_{1}\left(\sum_{t=0}^{T-1}\gamma^{kT+t}r_{(k,t)}^{\mathrm{CM\_trad}}\right)+\kappa_{2}\gamma^{k}\xi_{k+1}.$ (49)
    Note that $\xi_{k+1}=A_{\pi^{\text{sup}}}(S_{k+1}^{\text{sup}},a_{k+1})|_{a_{k+1}^{\text{inf}}\sim\pi^{\text{inf}}(\cdot|S_{k+1}^{\text{inf}})}$ represents the amount of control performance loss that the information recipient will suffer from, as a consequence of its control agent having to make a decision at the next control interval $k+1$ based on the partial/imperfect observation $S_{k+1}^{\text{inf}}$ according to the inferior policy $\pi^{\text{inf}}(\cdot|S_{k+1}^{\text{inf}})$ instead of based on the full/perfect observation $S_{k+1}^{\text{sup}}$ according to the superior policy $\pi^{\text{sup}}(\cdot|S_{k+1}^{\text{sup}})$. The partial/imperfect observation received at control interval $k+1$ is affected by the “When and What to Communicate” decisions made at the beginning of control interval $k$ by the communication agent for the information source. When $\xi_{k+1}$ in (49) corresponds to the IVoMI of a piece of information, we can determine whether the instance of information generated at control interval $k$ will be transmitted or not by the information source. When $\xi_{k+1}$ corresponds to the IVoII of a piece of information, we can determine what processed information should be shared by the information source.
For the “How to Communicate” Problem, we can formulate an SSDP to optimize the communication decisions made at each communication interval. Specifically, the reward function of the SSDP can be written as
@@
Note that the $r_{(k,t)}^{\mathrm{CM\_trad}}$ component is present in the reward of every communication interval, while the IVoII component, i.e., $\xi_{k+1}$, is only present in the last communication interval of each control interval when $t=T-1$. The imperfect observation received by the control agent at control interval $k+1$ is affected by the degree of communication nonideality the transmitted data has experienced, which is in part a result of the sequence of decisions on communication actions $a_{(k,0)}^{\text{CM}},a_{(k,1)}^{\text{CM}},\ldots,a_{(k,T-1)}^{\text{CM}}$ that are made during control interval $k$ by the communication agent.
@@
Case 10.
The state space, action space, and reward function of the SSDPs for making communication decisions in Case 9 are given below. The SSDPs can be solved by a centralized communication agent at the BS; or by the decentralized communication agents at each predecessor $i-1\in\mathcal{L}$. The algorithms to solve the SSDP models are out of the scope of this paper. We refer the readers to [27] for more details on a similar example.
@@
    • When and What to Communicate:
      $S^{\rm CM}_{(k)}=\{\bm{G}_{(k,0)},\{\tau_{i,k}\}_{i-1\in\mathcal{L}},\{u_{i,k^{\prime}}\}_{k^{\prime}=k-\tau_{\mathrm{max}}+1}^{k}\}$
@@
      $S^{\rm CM}_{(k,t)}=\{\bm{G}_{(k,t)},q^{\rm CAM}_{(k,t)},\{u_{i,k^{\prime}}\}_{k^{\prime}=k-\tau_{\mathrm{max}}+1}^{k},t\}$,
@@
    • When and What to Communicate:
      $a^{\rm CM}=\{\phi_{i-1,k}\}_{m\in\mathcal{M},i-1\in\mathcal{L}}$
@@
    • When and What to Communicate:
      $r^{\rm CM}_{(k,t)}$ $=\kappa_{1}\left(\sum_{m=0}^{M-1}\sum_{t=0}^{T-1}\gamma^{kT+t}r_{m,(k,t)}\right)$
      $+\kappa_{2}\sum_{i-1\in\mathcal{L}}\gamma^{k}\xi_{i,k+1}.$ (53)
@@
      $\left\{\begin{array}[]{ll}\kappa_{1}\sum_{m=0}^{M-1}r_{m,(k,t)},&\ 0\leq t<T-1\\ \kappa_{1}\sum_{m=0}^{M-1}r_{m,(k,t)}+&\\ \kappa_{2}\sum_{i-1\in\mathcal{L}}\xi_{i,k+1}^{\rm I},&\ t=T-1\\ \end{array}\right..$ (57)
@@ VIII Conclusion and Future Work
The significance of this paper is to establish a canonical framework for modeling the value of the information shared over 6G V2X networks, specifically for vehicle control decisions under various forms of uncertainty. To lay the theoretical groundwork, we have presented the SSDP model as an extension of the MDP model, explicitly accounting for useful information that may be missing from the state. This extension accommodates both fully observable and partially observable scenarios. Based on the SSDP model, we have proposed a mathematically rigorous and practical method to identify the set of information that can enhance decision-making if made available.
@@
While the proposed framework holds great promise for enabling the joint optimization of stochastic, sequential control and communication decisions across diverse networked control systems, several key challenges present compelling directions for future research.
Firstly, accurate VoI estimation by control agents is crucial for providing precise reward signals that guide the communication agents’ decision-making optimization. The estimation methods discussed in this paper, particularly those for IVoI, still have room for improvement. This is because IVoI is essentially an advantage function, and existing estimation methods for advantage functions are primarily focus on learning optimal policies rather than accurately estimating the advantage functions themselves. As a result, the precision of the learned values - especially for those states and actions that are infrequently visited under policies with full or perfect observations - may not be sufficient for accurate VoI estimation.
Secondly, the construction of communication state variables in SSDP models plays a critical role in the performance of the learned communication policies. Generally speaking, the communication state should comprise two components. The first component includes Channel State Information (CSI), Queue State Information (QSI), and other elements typically defined in MDP models for communication problems. Meanwhile, the second component pertains to IVoI, which theoretically should include any information that is useful in predicting the IVoI at the recipient. The IVoI, as an advantage function, depends on the recipient’s full or perfect observations, the action determined by the policy under partial or imperfect observations, and the neural network parameters for the critic of the policy with full or perfect observations. However, since the communication agent at the information source cannot obtain or construct the state using all this information, developing a systematic method for communication state construction that is independent of specific control tasks remains an open issue.
Lastly, since the VoI only appears in the reward of the last communication interval within each control interval in (10), the challenge of learning efficient communication policies under a sparse reward setting must be addressed.

### v2

@@ VII-C Optimization Objective
Case 10.
In Case 9, the optimization objective of the communication agents $i-1\in\mathcal{L}$ is to maximize the sum V2I throughtput while minimizing the performance degradation of the followers $i$ due to non-ideal communications.
@@
Similarly, let $G_{m,(k,t)}$ denote the instantaneous channel gain of V2I link $m$; $G_{i-1,B,m,(k,t)}$ the interference channel gain from V2V link $i-1$ transmitter at predecessor $i-1$ to V2I link $m$ receiver at the BS; $G_{B,i-1,m,(k,t)}$ the interference channel gain from V2I link $m$ transmitter at a vehicle to V2V link $i-1$ receiver at follower $i$; and $G_{j-1,i-1,m,(k,t)}$ the interference channel gain from V2V link $j-1$ transmitter at predecessor $j-1$ to V2V link $i-1$ receiver at follower $i$ over the sub-channel $m$.
@@
$\gamma_{m,(k,t)}=$
$\frac{P^{\rm I}_{m}G_{m,(k,t)}}{\sigma^{2}+\sum\limits_{i-1\in\mathcal{L}}\phi_{i-1,k}\theta_{i-1,m,(k,t)}P^{\rm V}_{i-1,m,(k,t)}G_{i-1,B,m,(k,t)}},$ (37)
@@
respectively. $\sigma^{2}$ is the power of channel noise which satisfies the independent Gaussian distribution with a zero mean value. $I_{i-1,m,(k,t)}$ in (38) is the total interference power received by V2V link $i-1$ over sub-channel $m$, where
@@
Each predecessor $i-1\in\mathcal{L}$ maintains a buffer to store its CAMs. Let $q^{\rm CAM}_{i-1,(k,t)}$ denote the queue length, i.e., the number of CAMs in the buffer of predecessor $i-1$ at communication interval $(k,t)$. At the beginning of the first communication interval $(k,0)$ of each control interval $k\in\mathcal{K}$, predecessor $i-1$ samples the driving status for control interval $k$ and generates a CAM. If the communication agent decides to transmit this CAM ($\phi_{i-1,k}=1$), it is stored in the buffer. During each communication interval $(k,t)$, where $t\in\mathcal{T}$, the queue length $q^{\rm CAM}_{i-1,(k,t)}$ is decreased by $\Delta t\times r^{\rm CAM}_{i-1,(k,t)}$, which represents the number of CAMs transmitted during this communication interval.
Since any CAM not fully transmitted by the end of the last communication interval $(k,T-1)$ is discarded, the queue is reset at the start of each new control interval $k+1$. Specifically, at communication interval $(k+1,0)$, $q^{\rm CAM}_{i-1,(k+1,0)}$ is set to $1$ if a CAM is scheduled for transmission ($\phi_{i-1,k+1}=1$), or 00 otherwise ($\phi_{i-1,k+1}=0$). Because the queue length changes abruptly at the boundaries between control intervals, we use $q^{\rm CAM}_{i-1,(k,T)}$ to denote the queue length at the end of control interval $k$, i.e., after the final communication interval $(k,T-1)$. Note that $q^{\rm CAM}_{i-1,(k,T)}\neq q^{\rm CAM}_{i,(k+1,0)}$ due to this discontinuity.
@@
$\left\{\begin{array}[]{ll}\phi_{i-1,k},&\mathrm{if}\quad t=0\\ \max\left[0,q^{\rm CAM}_{i-1,(k,t-1)}-r^{\rm CAM}_{i-1,(k,t-1)}\Delta t\right],&\mathrm{if}\quad 0<t\leq T\\ \end{array}\right..$ (44)
For each follower $i$, if a CAM is fully delivered by the end of control interval $k$, the receiver will update its stored CAM with the newly received one at the beginning of the next control interval $k+1$ and uses the updated information for vehicle control. If a CAM is not received or only partially received, the receiver has two option: (1) replace the missing CAM with dummy bits (e.g., zeros); (2) continue using the last fully received CAM.
With option (2), when the most recent available CAM of follower $i$ at the beginning of control interval $k$ corresponds to the CAM sampled at the predecessor $i-1$ in control interval $k-\tau_{i,k}$, we refer to $\tau_{i,k}$ as the observation delay at V2V receiver $i$, which is derived from the queue length as
@@
where $\gamma$ is the communication reward discount factor. The first term corresponds to the discounted sum throughput of all V2I links $m\in\mathcal{M}$ over all the communication intervals of a control episode, and the second term corresponds to the EVoMI $\Xi_{i}^{\rm M}$ or EVoII $\Xi_{i}^{\rm I}$ of follower $i$ defined in Case 6. The weight factors $\kappa_{1}$ and $\kappa_{2}$ indicate the relative importance of minimizing control performance degradation versus maximizing the V2I throughput.
@@
where the expectation is taken with respect to the probability distribution of the state-action trajectories when the communication policy is $\pi^{\rm CM}$ and vehicle control policy is $\pi^{\mathrm{inf}}$.
@@ VII-D VoI-oriented Optimization for Vehicular Communication
When addressing the problem of “When and What to Communicate”, different approaches can be adopted depending on the decision-making time scale:
  • Across Multiple Control Intervals: When decisions are made statically or semi-statically across multiple control intervals, one can evaluate $J^{\mathrm{CM}}$ under different communication decisions and choose the decision that maximizes $J^{\mathrm{CM}}$. To this end, the EVoI can be estimated as outlined in Section V, and $J^{\mathrm{CM\_trad}}$ can be evaluated using analytical tools such as queuing theory or through simulation.
  • Every Control Interval: When decisions are made dynamically at each control interval, an SSDP can be formulated to optimize communication decisions at each step. In this case, the reward function of the SSDP can be expressed as
    $r^{\rm CM}_{(k,t)}=\kappa_{1}\left(\sum_{t=0}^{T-1}\gamma^{t}r_{(k,t)}^{\mathrm{CM\_trad}}\right)+\kappa_{2}\xi_{k+1},$ (50)
    where $\xi_{k+1}=A_{\pi^{\text{sup}}}(S_{k+1}^{\text{sup}},a_{k+1})|_{a_{k+1}^{\text{inf}}\sim\pi^{\text{inf}}(\cdot|S_{k+1}^{\text{inf}})}$ is the IVoI. This terms captures the expected control performance degradation the recipient may experience due to making a decision at the next control interval $k+1$ based on the partial/imperfect observation $S_{k+1}^{\text{inf}}$ rather than the full/perfect observation $S_{k+1}^{\text{sup}}$.
    The partial/imperfect observation available at control interval $k+1$ is affected by the “When and What to Communicate” decisions made by the communication agent at control interval $k$. Specifically,
    • If $\xi_{k+1}$ corresponds to the IVoMI, it helps determine whether the raw information generated at control interval $k$ should be transmitted.
    • If $\xi_{k+1}$ corresponds to the IVoII, it guides the selection of which processed information should be shared.
For the “How to Communicate” problem, an SSDP can be formulated to optimize the communication decisions made at each communication interval. Specifically, the reward function of the SSDP can be expressed as
@@
Note that the $r_{(k,t)}^{\mathrm{CM\_trad}}$ component is present in the reward of every communication interval, while the IVoII component, i.e., $\xi_{k+1}$, is only present in the last communication interval of each control interval when $t=T-1$. The imperfect observation received by the control agent at control interval $k+1$ is affected by the degree of communication nonideality the transmitted data has experienced, which is in part a result of the sequence of decisions on communication actions $a_{(k,0)}^{\text{CM}},a_{(k,1)}^{\text{CM}},\ldots,a_{(k,T-1)}^{\text{CM}}$ that are selected during control interval $k$ by the communication agent.
@@
Figure 5: Performance comparison between the VoI-oriented optimization approach and a baseline method that always transmits the predecessor’s acceleration, evaluated in a simplified C-V2X network supporting vehicle-following control applications.
Case 11.
The state space, action space, and reward function of the SSDPs for making communication decisions in Case 10 are given below. The SSDPs can be solved by a centralized communication agent at the BS; or by the decentralized communication agents at each predecessor $i-1\in\mathcal{L}$. The algorithms to solve the SSDP models are out of the scope of this paper. We refer the readers to [34] for more details on a similar example.
@@
    • When to Communicate:
      $S^{\rm CM}_{(k)}=\{\bm{G}_{(k,0)},u_{i-1,k}\}$
@@
      $S^{\rm CM}_{(k,t)}=\{\bm{G}_{(k,t)},q^{\rm CAM}_{(k,t)},\{u_{i-1,k^{\prime}}\}_{k^{\prime}=k-\tau_{\mathrm{max}}+1}^{k},t\}$,
@@
    • When to Communicate:
      $a^{\rm CM}=\{\phi_{i-1,k}\}_{i-1\in\mathcal{L}}$
@@
    • When to Communicate:
      $r^{\rm CM}_{(k,t)}$ $=\kappa_{1}\left(\sum_{m=0}^{M-1}\sum_{t=0}^{T-1}\gamma^{t}r_{m,(k,t)}\right)$
      $+\kappa_{2}\sum_{i-1\in\mathcal{L}}\xi_{i,k+1}.$ (54)
@@
      $\left\{\begin{array}[]{ll}\kappa_{1}\sum_{m=0}^{M-1}r_{m,(k,t)},&\ 0\leq t<T-1\\ \kappa_{1}\sum_{m=0}^{M-1}r_{m,(k,t)}+&\\ \kappa_{2}\sum_{i-1\in\mathcal{L}}\xi_{i,k+1},&\ t=T-1\\ \end{array}\right..$ (58)
To demonstrate the advantages of VoI-oriented optimization, we conduct experiments on a simplified C-V2X network supporting vehicle-following control applications, as considered in Case 9 and Case 10, with $M=1$ V2I link and $L=1$ V2V link. Focusing on the “When to Communicate” problem, we consider that if a CAM is scheduled for transmission, the predecessor always transmits at maximum power.
Fig. 5 shows a performance comparison between the VoI-oriented optimization approach and a baseline method that always transmits the predecessor’s acceleration. The VoI-oriented approach selectively suppresses CAM transmissions when the predecessor’s acceleration is zero, recognizing that such information holds no value for the follower, as indicated by the IVoI shown in Fig. 5. As a result, the V2I link experiences less interference from the V2V link during these periods, resulting in higher throughput compared to the baseline. Meanwhile, the position error ($e_{pi,k}$), velocity error ($e_{vi,k}$), and follower’s acceleration trajectories ($acc_{i,k}$) remain nearly identical under both methods, demonstrating that omitting the transmission of zero acceleration has a negligible impact on control performance.
@@ VIII Conclusion and Future Work
The significance of this paper is to establish a canonical framework for modeling the value of the information shared over 6G V2X networks, specifically for AD decisions under various forms of uncertainty. To lay the theoretical groundwork, we have presented the SSDP model as an extension of the MDP model, explicitly accounting for useful information that may be missing from the state. This extension accommodates both fully observable and partially observable scenarios. Based on the SSDP model, we have proposed a mathematically rigorous and practical method to identify the set of information that can enhance decision-making if made available.
@@
Although the framework has been developed with CAV systems in mind, it holds great promise for enabling the joint optimization of stochastic, sequential control and communication decisions across diverse networked control systems (NCS) and autonomous/artificial Internet of Things (AIoT) applications [12], including networked robots, smart grid, industrial automation systems, unmanned aerial vehicle (UAV) swarms, and more. Several key challenges present compelling directions for future research.
Firstly, accurate VoI estimation by control agents is crucial for providing precise reward signals that guide the optimization of communication agents’ decision-making. The estimation methods discussed in this paper, particularly those for IVoI, still leave room for improvement. This is because IVoI is essentially an advantage function, and existing estimation methods for advantage functions are primarily designed to support policy learning, rather than to achieve high-precision value estimates. As a result, the precision of the learned values - especially for those states and actions that are infrequently visited under policies with full or perfect observations - may be insufficient for accurate VoI estimation. One potential remedy is to combine Method A with Method B or Method C, aiming to achieve high estimation accuracy with a reasonable amount of experience.
Secondly, the construction of communication state variables in SSDP models plays a critical role in the performance of the learned communication policies. Generally speaking, the communication state should comprise two key components. The first includes Channel State Information (CSI), Queue State Information (QSI), and other elements typically defined in MDP models for communication problems. The second component pertains to IVoI, which theoretically encompasses any information that is useful in predicting the IVoI at the recipient. The IVoI, as an advantage function, depends on the recipient’s full or perfect observations, the action determined by the policy under partial or imperfect observations, and the neural network parameters for the critic of the full/perfect- observations policy.
Ideally, the communication states should incorporate all of the above information. However, in practice, the composition of communication states varies across tasks, depending on what information is available to the communication agents. Developing a systematic method for constructing communication state representations that is agnostic to specific control tasks remains an open issue.
Lastly, VoI appears only in the reward of the last communication interval within each control interval in (11), resulting in a sparse reward problem. This issue is further exacerbated in multi-agent RL settings, which are common in RRM problems, due to increased complexity and coordination challenges among agents. Therefore, learning efficient communication policies under sparse reward conditions remains a critical issue to address.